Talamoto Co. manufactures a single product that goes through two processes Mixing Department for September. mixing and cookinMultiple $1,094,800. $957,600. $932,000. $856,000. $887,000.
Answer #1

Unit transferred out = 28000+72000-16000 = 84000

Equivalent unit of material P = 84000+16000 = 100000

Equivalent unit of material Q = 100000

Equivalent unit of conversion = 84000+(16000*50%) = 92000

Cost per equivalent unit of material P = (120000+180000)/100000 = 3

Cost per equivalent unit of material q = (110000+165000)/100000 = 2.75

Cost per equivalent unit of conversion = (165000+354800)/92000 = 5.65

Cost of unit transferred out = (3+2.75+5.65)*84000 = 957600

So answer is b) $957600